OpenStudy (anonymous):

please help!! If θ is an angle in standard positionand B(-3,4) is a point on the terminal side of the angle, what is the value of sin θ ?

OpenStudy (debbieg):

Do you know the identities for the trig functions, in terms of the coordinates of a point on the terminal side? You'll want to look them up and commit them to memory. :) If (x,y) is on the terminal side of θ, then: sinθ=y/r where \(r=\sqrt{x^2+y^2}\) (the distance of the point (x,y) from the origin.

OpenStudy (anonymous):

thank you so much! i remember how to do it now!